Suppose that you plan to borrow $20,000 student loans to attend UM-Dearborn. You are

considering borrowing the loan from SallieMae. SallieMae offers two options for the repayment

of your loan. One is the deferred repayment option and the other is interest repayment option. The

APR for the deferred repayment option is 6.75% and the APR for the interest repayment option is

5.75%. You plan to finish your undergraduate study in UM-Dearborn within five years. The two

repayment options are described as below:

Deferred repayment option: You make no scheduled student loan payments for 5 years while you

are in school and in 1 year of the grace period after you graduate. However, the unpaid interest

every month will be added to your principal amount at the end of your grace period. After the

grace period, the total amount your will pay will be equal to the principal you borrow and the

accumulated unpaid interest. Each month you will be required to pay the same amount, which

includes interest and the required principal repayment. You are required to completely pay off

your loan within 10 years.

Interest repayment option: You pay interest every month when you are in school and in grace

period. After the grace period, you will start to pay the principal of the loan. Each month you will

be required to pay the same amount, which includes interest and the required principal repayment.

You are required to completely pay off your loan within 10 years.

1. Please use Excel to work on the following questions:

1) Set up the loan amortization tables for the loans with these two different repayment

options, respectively.

2) How much interest will you pay in total when you pay off your loan offered by these two

different repayment options, respectively?

3) How much will you pay in total, including interest and principal, for these two different

repayment options, respectively?
Answer #1

The entire question is answered on the excel along with detailed formulas. Below is the detailed explaination of the solution

Deferred Repayment Option Formulas Interest Repayment Option Formulas
Borrowed Amount                                            20,000.0 Given                                         20,000.0 Given
APR 6.75% Given 5.75% Given
No Principal Repayment till (years) 6 Given 6 Given
Interest paid/ accumulated till 6 years                                              8,100.0 =C4*C3*C5                                            6,900.0 =E4*E3*E5
Effective Principal after 6 years (accumulated unpaid interest)                                            28,100.0 =C3+C6                                         20,000.0 =E3
Repayment of entire loan in (years) 10 Given 10 Given
Repayment of entire loan in (months) 120 =C8*12 120 =E8*12
Monthly Rate (%) 0.6% =C4/12 0.5% =E4/12
EMI/month 322.66 =PMT(C10,C9,-C7,0,0) 219.54 =PMT(E10,E9,-E7,0,0)
Total Repayment in 10 years                                            38,718.7 =C11*120                                         33,244.6 =(E11*120)+E6
Total Principal Repayment                                            20,000.0 =C3                                         20,000.0 =E3
Total Interest paid                                            18,718.7 =C12-C13                                         13,244.6 =E12-E13